### v2

# Deciphering the global production network from cross-border firm transactions

Abstract
Critical for policy-making and business operations, the study of global supply chains has been severely hampered by a lack of detailed data. Here we harness global firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products. Transforming this data to a directed network, we find that products are clustered into three large groups including textiles, chemicals and food, and machinery and metals. European industrial nations and China dominate critical intermediate products in the network such as metals, common components and tools, while industrial complexity is correlated with embeddedness in densely connected supply chains. To validate the network, we find structural similarities with two alternative product networks, one generated via LLM queries and the other derived by NAFTA to track product origins. We further detect linkages between products identified in manually mapped single sector supply chains, including electric vehicle batteries and semi-conductors. Finally, metrics derived from network structure capturing both forward and backward linkages are able to predict country-product diversification patterns with high accuracy.
Keywords: Supply chains, resilience, policy, network science
Disclaimer: This paper contains the views of the authors and not necessarily those of the UK Government or its Ministers.
Acknowledgements: We would like to thank Christian Diem and Tobias Reisch for their useful input and comments, and Shaun Hoang for his research assistance. The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP). NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme.
Introduction
As acutely highlighted by the global pandemic, limited insight into global supply chains can have serious and immediate consequences. However, until quite recently, supply chain analysis was limited to specific sectors or highly aggregate economy-wide transaction data which lacks detail and hence is not very useful or sufficiently accurate for many policy and research questions. Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product level supply chain network.
Traditionally, flows and transactions between sectors in the economy have been captured via input-output tables (Leontief, 1986). These capture monetary flows between firms within defined industries, and are typically produced by national statistical agencies. Some countries, such as the US and Mexico, produce these at quite a detailed level (4 digit NAICS - or approx 400 industries). However, in most countries these are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors). And, importantly, they don’t capture cross border flows limiting their potential for global supply chain analysis. There have been a handful of efforts to build integrated input-output networks that include inter-country flows, but again these are extremely aggregate in nature. For example, the World Input-Output Dataset (WIOD) (Timmer et al., 2012) has just 56 sectors and 40 countries (2010-14) or the OECD World Input-Output Table (WIOT) with 36 sectors and 69 countries (1995-2020). A variety of studies have investigated the structure of the World Input-Output Network (WION), finding strong geographical clustering of sectors within neighbouring countries, but also some instances of strong inter-country clustering in sectors such as textiles and transport equipment Piccardi et al. (2018).
In recent years there has been a huge push to better collect and make available for research firm level transaction data (Pichler et al., 2023). Research using this data has indicated that, for example, the propagation of shocks at the firm level has distinct characteristics and magnitudes which cannot be approximated by aggregate data (Diem et al., 2022, 2024; Inoue and Todo, 2024). Up until very recently, however, most firm level data was collected by single countries (e.g. VAT data) or for a specific purpose (i.e., shipping data), and hence had very limited coverage.
At the other end of the spectrum are efforts to manually reconstruct detailed supply chains for specific sectors. These studies collect data from a variety of sources including internal systems like procurement, inventory management, and logistics, as well as external sources such as supplier data and transportation information, to build a detailed picture of the key inputs needed to produce often strategically important goods. For example, there have been careful efforts to map out antibiotic (Klimek et al., 2023), semi-conductor (Haramboure et al., 2023) and battery supply chains (Caravella et al., 2024). However, these supply chain mapping exercises tend to focus on material inputs - critical minerals and chemicals - rather than machinery, tools and other necessary inputs such as cleaning equipment. And, critically, their fragmented nature makes them not suited to many questions which require a wider structural lens on product inter-connectivity.
A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks. Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters, sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad based shocks such as the pandemic or financial crisis. Much of this literature aims to quantify node importance - whether it be nodes that influence the degree of spreading processes, or those that are most susceptible to impacts. A number of papers suggest that the US IO network follows a power law structure (Acemoglu et al., 2012; Han and Goetz, 2019; Chopra and Khanna, 2015), and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades. Blöchl et al. (2011) proposed two centrality measures to capture the impact of shocks: random walk centrality to capture most immediate impacts, and betweenness to capture where a shock lingers longest. The authors apply these metrics to the OECD WIOT, finding ’wholesale and retail trade’ is most often the node with highest random walk centrality while ’motor vehicles’ have the highest betweenness.
Supply chain linkages are also key to longer term impacts such as diversification of countries and regions into new economic activities, otherwise referred to as ’upgrading’ (Kaplinsky, 2000). Originally proposed as a mechanism to aid countries with natural resources diversify into productive activities (Hirschman, 1958; Gereffi et al., 1999), it is widely understood today that embeddedness in global supply chains aids both domestic productivity and export diversification (Pahl and Timmer, 2020). While a wide array of mechanisms have been proposed, many focus on accumulation of technological capabilities and the diffusion of knowhow, as espoused by the Evolutionary Economic Geography (Nelson and Winter, 2002; Boschma and Frenken, 2006) and Innovation Systems (Edquist, 2013) literatures. From this perspective, supply chain linkages are thought to drive knowledge diffusion via, for example, demands for quality upgrading, worker flows and increased competition (Blomström and Kokko, 1998; Criscuolo et al., 2017). These effects are thought to be particularly strong in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals (Javorcik, 2004), but there are greater uncertainties about the potential for such linkages to induce upgrading when it comes to linkages spanning large geographical and technological distances (Pietrobelli and Rabellotti, 2011).
The empirical evidence on the benefits of participation in global value chains is mixed (as reviewed by Kowalski et al. (2015)), finding a strong influence on manufacturing productivity and export sophistication, but mixed impact, for example, on domestic employment (Pahl and Timmer, 2020; Gonzalez, 2016). In terms of direction, both backward linkages (i.e., diversification into more complex downstream goods) and forward linkages (i.e., into input or supplier goods) have been proposed as channels for diversification. Methodologically closest to our work, Bahar et al. (2019) find evidence for the role of forward linkages in international export product diversification patterns over a 10 year period. Linkages in this case, however, were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
Beyond the short-medium term impacts of shocks, a recent literature studies the influence of production network structure on local economic growth and resilience. Taking an evolutionary stance, in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms (Boschma, 2015). Empirical evidence at a subnational level supports the idea that regions’ diversification patterns are conditioned on their ’role’ in national production networks. For example, US cities tend to enter new industries that are ’related’ to those that already exist (Essletzbichler, 2017), where relatedness is measured in terms of supply chain proximity (Fan and Lang, 2000). When it comes to resilience, Han and Goetz (2019) show that weighted centrality metrics derived from the US IO network are associated with employment resilience for US counties during the 2007-2009 recession, while Cainelli et al. (2019) find that EU regions with highly integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
Here we aim to infer detailed ’product recipes’, or the set of unique inputs needed in the production of a specific product. In other words, we aim to capture the input materials and tools used in product that are specific to a particular product, and not common across many products (e.g., packaging). Distinct from two recent efforts build a similar product network using firm-level transaction data from Italy (Fessina et al., 2025) and Japan (Inoue and Todo, 2024), we aim to capture global supply chains (as opposed to local supply chains) via global firm purchasing patterns. To do this, we harness ’next generation’ global firm-level transaction data via a UK government programme, GSCIP. We use an aggregated subset of data from GSCIP for the years 2021-2023, prior to additional augmentation of the transactional data completed in 2025. The subset of data that we use is orders of magnitude larger than single country datasets typically in use today, and covers 20m global firms, and 1 billion cross-border transactions. The data was assembled from a variety of sources, from customs to firm-owned data, and matched at a firm level.
To detect the set of inputs required to make a specific product, we compute the excess purchase of inputs (products) for the set of firms that make a specific product relative to how often an average firm purchases the input. This latter step effectively down-weights common or ubiquitous inputs. This approach echoes that of O’Clery et al. (2021), in which the authors inferred country-product diversification paths by extracting the set of common products produced in the set of countries in which a specific new product appeared later. In this case, one can think of this approach as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate each set of input linkages for each product. In this way, we capture both inputs or ingredients, and complex tools and packaging.
These relationships can be represented and analysed as a network in which nodes are products, and directed weighted edges represent the relative importance of the input. Analysis of the meso-structure of this network suggests that textiles’ supply chains are largely isolated from wider manufacturing, while chemicals are typically more inter-connected with food supply chains than machinery and metals. European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of Hidalgo and Hausmann (2009)) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
While there is no ‘ground truth’ data available, we have undertaken significant efforts to validate the network structure. We found structural similarities with an alternative AI-generated product network (Fetzer et al., 2024), and a product supply chain dataset created by NAFTA to track product origins (Conconi et al., 2018). In addition, we found dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries and semi-conductors. Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy, finding evidence for both forward and backward linkages.
Results
We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as shipping data. These sources have been integrated using AI-based tools to, for example, disambiguate firm names and match them. The data we used represents a static cut which was made available to the research team via a UK government programme (GSCIP), includes firm level attributes for over 20m firms and 1 billion inter-firm transactions between 2021-23. All transactions are between firms located in different countries. The GSCIP data contains additional firm-level information from other sources, such as firm ownership structures.
We have a number of firm and transaction attributes, a selection of which is shown in Figure 1 B, including dollar transactions for specific products between firm pairs. As shown in Figure 1 C-D, the data has high coverage across developed economies, and those with high trade (exports plus imports) volumes such as China and Vietnam. It also has good coverage in some middle income countries such as India and Peru, but with lower coverage overall in less economically active regions such as Africa.
Figure 1: We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. [A] Firm level transaction data has been compiled from a variety of sources. [B] This static cut of GSCIP data includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions. [C-D] The data has very good coverage of firms in developed economies, with lower coverage in less economically active regions such as Africa. Top 50 countries by firm count shown.
We infer the importance of each product as an input to any other product from the firm transaction data. We aim to derive a metric to capture the ’importance’ of product $i$ for the production of product $j$. In essence, we compute the ’excess’ purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example in Figure 2 B, we show that while aluminium panels are purchased very frequently by washing machine firms (relative to all firms), cardboard boxes are often purchased by all types of firms including washing machine firms and hence they are a very common input (and are thus down-weighted in our scheme).
More precisely, we set stringent definitions to identify firms that both produce and purchase specific products:
  • A firm $k$ produces product $j$ if they sell a larger monetary value of $j$ than the average firm that sells product $j$.
  • A firm $k$ buys product $i$ if they purchase a larger monetary value of $i$ than the average firm that buys product $i$.
The set of all firms is denoted $S$. For each product $j$, we form a subset of firms that produce product $j$, denoted $S_{j}$. A subset of these buy product $i$, $S^{i}_{j}$. Finally, $S^{i}$ is the set of firms that purchase product $i$ (irrespective of what they produce).
Our ratio is the share of firms who produce $j$ that also buy $i$ divided by the overall share of firms that buy $i$:
$A_{i,j}=\frac{|S^{i}_{j}|/|S_{j}|}{|S^{i}|/|S|}$
If $A_{i,j}>1$, then product $i$ is more prevalent in the purchasing basket of firms that produce $j$ than the wider set of all firms. We can also set a minimum for the size of the set $|S^{i}_{j}|$ (the number of firms that both produce $j$ and buy $i$), which we refer to as parameter ’firmcount’. In essence, for low values of firmcount, we have few ’pieces of evidence’ for the edge weight (and so for some applications we might want to set a higher firmcount threshold, which would make our network sparser). In the analysis below, we often show results for a range of values of both the edge weight threshold and firmcount.
As shown in Figure 2 B, we pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
Figure 2: We infer the importance of each product as an input to any other product from the firm transaction data. [A] We compute the excess purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example here we show that while aluminium panels are purchased very frequently by washing machine firms relative to all firms, cardboard boxes are purchases by all types of firms including washing machine firms and hence they are a very ubiquitous input (and are thus down-weighted in our scheme). [B] We pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
We transform the product input relations into a network, and find three main clusters representing machinery and metals, chemicals and food, and textiles. We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$. In Figure 3 A, we observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold has no major effect. In practice, the majority of the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$. In many cases, and also in the SI, we also show results when varying these parameters.
This network can be visualised using a ’spring’ based visualisation algorithm (Force Atlas in Gephi was used here), as shown in Figure 3 B. We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors. However, rather than rely on an external classification which bears no relation to the inter-connectivity we observe in our data, we use a multi-scale community detection algorithm to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’. These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network. These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation (Fortunato, 2010).
While there are many distinct approaches to community detection suited to different applications (Fortunato, 2010), in our case, we employ a multi-scale algorithm based on random walkers on a network (Delvenne et al., 2010). Essentially, as the walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, we detect larger and larger clusters with strong internal connectivity. Hence, as ’time’ goes on, the algorithm detects fewer but larger communities. We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters, the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity. In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
Here, shown in Figure 3 C-E, we show three network partitions into communities at three distinct ’times’ (see SI for more details and further analysis). The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown. We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves). We observe that over time, food merges with plastics and then chemicals, and both machinery clusters merge. Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged signifying their isolation. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
Figure 3: We transform the product input relations into a network. [A] The product supply chain network can represented by an adjacency matrix in which a value in entry $A_{i,j}$ denotes a non-zero edge weight between node (product) $i$ and node $j$. We observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firm-count parameter, the edges become more sparse, but the edge weight threshold has no major effect. [B] The network can also be visualised using a visualisation algorithm (Force Atlas was used here). [C-D] We use a multi-scale community detection algorithm to uncover the ’meso’ structure of the network. As the algorithm effectively ’forces’ communities to merge into larger clusters, we find that textiles is highly isolated in the sense that it remains unmerged over many scales. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. In the SI we show that machinery products including those related to cars and motors tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to be inputs to many products. Further, consistent with what we might expect, deploying a UN classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients.
Betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross. Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) paths transits that node. In Figure 4 A, we highlight nodes with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders. Nations which produce these key inputs have an important role in global supply chains, controlling so-called ’choke points’. To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) of 113 countries from global trade data (Comtrade 2024). In Figure 4 B, we show a global map with countries shaded by their share of top-50 BC products exported with RCA. We find European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
Highly industrialised nations engage in complex production processes via embeddedness global supply chains. To capture production complexity, we adopt an approach proposed by Hidalgo and Hausmann (2009) to infer product complexity (PCI) from patterns of geographical production. To measure the complexity of the supply chain in the vicinity of a product in our network, we compute the hub score of each node. This is similar to eigencentrality in the sense that it captures ’neighbour of neighbour’ linkages, but is suitable for directed networks such as ours. We find that the PCI (obtained for 2023 from the Harvard Atlas of Economic Complexity) has a rank correlation of 0.46 with node hub score. In essence, complex products are embedded in complex supply chains. At the country level, the country-level Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score. Hence, countries with high industrial complexity tend to export products which are highly embedded in densely connected global supply chains.
Figure 4: European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. [A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ have high betweenness centrality. [B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products. [C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures ’neighbour of neighbour’ linkages. [D] The Product Complexity Index (rank) correlates 0.46 with hub score. [E] Economic Complexity Index (ECI) has a high rank correlation of 0.72 with mean hub score. Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
The supply chain network correlates with other efforts to reconstruct inter-product linkages. First, we consider a recent effect to construct a similar type of product network using LLM queries (Fetzer et al., 2024). In essence, the authors asked an LLM what are the top input products for each product, and constructed a network based on its answers in a similar fashion to above.
Figure 5 A shows that the (binary unweighted) AI network is much sparser than our network (see adjacency matrix in inset). Its clear from the distribution of white nodes (those with zero in-degree) that the AI network tends to have much fewer edges in sectors such as chemicals, plastics and food. Overall, our network (the SC network) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community. We compare communities for the AI and SC network in Figure 5 C. We find that the AI and SC product networks have similar meso-scale structure, with the exception of much larger and stronger chemicals and food clusters in the SC network.
Next, we compute a simple correlation of the full edge set of AI network and SC network, finding they are correlated up to approx 0.35 for low values of firmcount (i.e., including uncertain edges). The in- and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree. In general, these correlations peak for low firmcount in the range 2-4 - i.e., we keep weak edges for a dense network. Hence, while the networks look visually very different, they display non-trivial similarity in both meso structure and in- and out-degree pattern.
We also find similarity to a NAFTA derived product-level IO network. We compare our SC network to a NAFTA derived product-level ’IO network’ from Conconi et al. (2018) which was constructed based on NAFTA ’Rules of Origin’ which aims to identify the origin country of goods in order to determine if tariffs apply. In this case we find a much lower edge correlation of 0.15, a binary in-degree correlation of 0.28 and out-degree correlation of 0.32. Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see Figure S3).
Figure 5: The product network correlates with other efforts to construct inter-product links or chains. [A] The ’AI network’ of Fetzer et al. (2024) is constructed via LLM queries. Nodes with zero in-degree are shown in white. [B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food. [C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in the SC network. [D-E] The correlation of the full edge set of AI network and SC network is approx 0.34, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
The product network picks up links identified in manually constructed single-sector supply chains. Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector specific supply chains. Here we take a selection of these, and investigate to what extent we pick up similar linkages using our method. We focus on two key sectors, electric vehicle batteries and semiconductors. For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 1 A/C. In other words, we show the edge pattern that we find using our method for the identified product sets. We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery.
We wish to show that statistically its very unlikely we could have found these links at random. To do this, we simulate 100k random networks with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random networks. The modularity score essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard null model which randomly allocates edges while keeping the node degree fixed). We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k subgraphs, finding an extremely low probability that our subgraph would be simulated at random (pvalue shown).
Figure 6: The product network picks up links identified in manually constructed single-sector supply chains. [A/D] We capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. [B/E] While products in electric vehicle battery are highly clustered in the centre, semi-conductors exhibit two distinct clusters composed of chemicals and machinery. [C/F] To show we pick up linkages between these products that are far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks. We observe that our subgraphs are in the very extremes of the distributions suggesting they would be extremely unlikely to occur at random.
Both downstream and upstream links predict country-product export diversification patterns. Finally, we probe whether metrics derived from our network are predictive of country-product diversification patterns. To do this, we use a similar approach to a large literature on density metrics in EEG (Hidalgo et al., 2018) to construct a ’downstream’ (and analogous ’upstream’) density metric. As illustrated in Figure 7 A/D, essentially, for a product $p$, this captures the share of top in-degree neighbouring products present in country $c$:
$d_{p,c}=\frac{\sum_{j\in J_{p}}I(A_{p,j})M_{p,c}}{\sum_{j\in J_{p}}I(A_{p,j})}$
where $J_{p}$ is the set of top $k=50$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0). We construct an equivalent ’upstream density’ metric.
Similar to O’Clery et al. (2021), we perform a standard Probit regression for the probability of a product appearance of the form:
$R_{p,c}=\Phi(\alpha+\beta_{d}d_{p,c}+\gamma_{p}+\eta_{c})$
where $R_{p,c}=1$ if product $p$ was present in country $c$ in 2021 but absent in 2016, $\Phi$ is a normal cumulative distribution function, and $\gamma_{p}$ and $\eta_{c}$ are product and country fixed effects respectively.
The tables in Figure B/E show that both downstream and upstream links predict country-product export diversification patterns with high predictive power (the AUC or ’area under the curve’ measures the model’s ability to distinguish between positive and negative cases) for both models with and without fixed effects. The AUC values we find are in line with the literature on predicting product entries in countries, and other similar efforts to predict diversification on a range of scales using density metrics derived from a variety of networks capturing ’capability overlap’ between products (Hidalgo et al., 2007; O’Clery et al., 2021).
Figure C/F shows, perhaps counter-intuitively but in line with some of the validation steps above, that the predictive power is highest for low firmcount. Hence, including ’weak links’ for which there is relatively lower certainty is informative for the model. The predictive power is also higher for lower entry threshold - i.e., we better predict small increases in Rpop (which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product) over a 5 year period compared to larger jumps.
Figure 7: Both downstream and upstream links predict country-product export diversification patterns. [A/D] Downstream/upstream-density captures a country’s presence in products that are downstream/upstream of the target product. [B/E] Both downstream and upstream links predict country-product export diversification patterns with fairly high predictive power (AUC) for cases with and without fixed effects. [C/F] The predictive power is highest for low firmcount (i.e., including ’weak links’), and for low rpop entry threshold (i.e., we predict small increases in rpop (product concentration relative to GDPPC).
Discussion
In this paper we develop a highly granular network tool to describe complex global inter-related product supply chains. We explore its broader structural characteristics, and investigate which countries dominate common input nodes with high betweenness, and which are active in highly embedded ’hub’ nodes. Despite an absence of ’ground truth’ data, we validate it in several ways including comparing it to a variety of other efforts to construct product level supply chain networks, and manually constructed data on sector specific supply chains.
There are a wide range of potential research and policy uses for this type of tool. From a policy viewpoint, this tool could be used to, for example, quickly map supply chains for new products which have not already been mapped manually, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs. More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains (Baldwin, 2006).
The network structure provides key insights on ’central nodes’ and clusters of highly interconnected products. Central nodes might be worthy of particular focus or protection. For example, the fact that steel has very high betweenness centrality indicates that it likely plays a role in many paths (chains) and might be seen as a critical good. This accords with the fact that the UK government moved aggressively to protect domestic production of domestic virgin steel in early 2025 in the face of impending closure of British Steel. More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks and policies such as tariffs in a more detailed, comprehensive manner than previously possible. In the case of product clusters, we can see network communities as a sort of alternative product hierarchy relative to official sectoral classifications. Highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be collectively targetted by policy interventions rather than individually.
In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods (Sato, 2014). Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions (Hertwich and Wood, 2018). This type of detailed network which captures - at least in theory - the full set of product relations could be used to track ’product life-cycle’ emissions as products are manufactured and then used in other manufacturing processes (Wang and Gupta, 2011).
Here we show that country-product export diversification is associated with both forward and backward linkages, with a slightly stronger signal for backward linkages. The network could be used as a policy tool to identify opportunities in the form of upstream and downstream products which could be targets for industrial policy support. This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors (Balland et al., 2019). Beyond nodes, it also sheds light on potential new edges. For example, it could be used to identify linkages which exist globally but not locally.
There are some limitations to our approach, as well as potential sources of noise. In order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years. This approach, however, does not currently allow for temporal analysis. Future work may investigate the extent to which disaggregating into distinct time periods and recomputing the network for each is feasible. Secondly, since our transactions are between entities located in different countries, we do not capture local linkages. For example, if the water used to produce beer is predominantly purchased locally, we may not capture it as a key input. Hence, we over-emphasise inputs and relationships in global value chains rather than local ones.
Third, there is no doubt some noise in our data due to the uneven composition of the firm-level data. For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of firms in our dataset, and some countries/sectors will be more represented than others. This issue is less problematic for our tool than say country-specific or firm-level analysis, but nonetheless it will introduce noise which we cannot fully mitigate against. Fourth, there may be other sources of noise, such as short term changes in purchasing patterns. For example, we noticed an unusual input to semiconductors - washing machines. But it turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips. Another source of noise is likely to be multi-product firms. While we have taken efforts to reduce this, by omitting firms that sell products in more than five HS Sections, and looking for common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
Fifth, there may be idiosyncratic differences between supply chain relationships within regions that we aggregate over by computing a network using all firm transactions. For example, battery supply chains may look different in the US vs Asia. While we have looked at some geographic variations here (see SI), in future work we may reconstruct the network using data for specific regions to further investigate this issue. Finally, we present the network at HS 4-digit level. It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and a more complicated validation exercise with fewer sources of external data to validate against. Again, we leave this to future work.
Methods
Processing of underlying firm data
The cut of data we used from GSCIP on firm-level transactions covers the time period between Jan 2021 and Dec 2023.
We take the following processing steps:
  • We assign all firms to an ’owner firm’ based on ownership information available to us (see SI).
  • We aggregate all firms to an ’owner-country’ level. In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
Construction of supply chain network
A concern about our approach might be that it could pick up spurious links due to the presence of firms in our dataset which make many products (and hence we cannot infer specific inputs to products from their purchases). Our methodology is designed to reduce this issue since we look at the purchases of not just one firm but all firms that make a specific product. To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of products such as Proctor and Gamble).
In order to further reduce noise, in addition to the firmcount parameter as mentioned above, we also drop any links from product $i$ to product $j$ in which the total value of the transactions of the firms in set $S^{i}_{j}$ is less than \$1000.
Network analysis
Community detection We use the Stability algorithm of Delvenne et al. (2010) to detect communities in the SC and AI networks. The algorithm detects partitions on a range of scales corresponding to random walker time on the network. The higher the resolution, the longer the time, the larger (and fewer) the communities. We use the ’full’ (non-linearised) specification for a directed Laplacian, with 100 iterations. In the SI we show the number of communities, variation of information across the iterations (a low value tells us we find similar partitions for different random seeds), and the stability score for our set of partitions corresponding to a range of timescales.
Modularity score We use the Modularity score to compare the connectivity of the subgraph induced by each of our sector-specific supply chains (EV batteries and semi-conductors) to 100k randomly generated subgraphs of the same size. Modularity (Newman, 2006) compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
Effectively, we decompose the directed Modularity score proposed by Leicht and Newman (2008) to evaluate a set of communities for a full network, and use it instead to evaluate the connectivity of a subgraph (i.e., a single community). This is mathematically coherent as Modularity is a sum over the connectivity within communities. Specifically, we compute:
$M_{G}=\frac{1}{m}\Large(\sum_{i,j\in G}X_{i,j}-\frac{Out_{i}In_{j}}{m}\Large)$
for node pairs $i,j$ in subgraph $G$, $X=A>0$ (unweighted adjacency matrix), $m$ is the number of edges in the full network, $Out_{i}$ is the out-degree of node $i$ and $In_{j}$ is the in-degree of node $j$.
Econometric analysis
To capture product entries, we wish to compute a measure of a countries’ concentration in a product. The traditional approach is Revealed Comparative Advantage (RCA) proposed by Balassa (1965) captures is the proportion of the country’s exports in a product divided by the global proportion of exports in a product. We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates (Hidalgo, 2021). Specifically, it compares a country’s per capita production levels in a product to the world’s overall per capita production of the product:
$Rpop_{p,c}=\frac{E_{p,c}/pop_{c}}{E_{p}/pop}$
where $E_{p,c}$ is the total exports of country $c$ in product $p$, $E_{p}$ is global exports of product $p$, $pop_{c}$ is the total population of country $c$ and poppop is the global total population.
We conduct the econometric analysis for a set of 235 countries. We drop products not in the main connected component of the network and those with relatively small global trade (less than 2 billion\$ in 2016) for a total of 683 products. Of a possible 235x683 observations, 86% were absent ($Rpop<0.05$) in 2016. Of these, 3.5% of products appeared by 2021 (with $Rpop>0.1$). While we keep the absence threshold fixed, we show results for a range of presence thresholds in Figure 7 C and F.
Supplementary Information
Overview of underlying data
Data sources and coverage
This paper relies on data provided by the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP). This contains firm-to-firm cross-border trade, derived from customs and shipping data, with details on the transacting firms compiled from tax filings, company websites, and other sources. Firm-to-firm transaction data is derived from administrative customs data. GSCIP amalgamates first-party customs data across a multitude of countries.
Our dataset includes transaction information for every country. While the dataset does not have first-hand customs data or another primary source of transaction information for every country, the data that we do have includes transactions between firms within the countries included and those out of sample providing partial coverage.
We are using a static cut of the data, which has further been restricted to the period from January 2021 to December 2023 as this period had the greatest coverage for the subset.
Data processing of underlying firm data
Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional processing of the data to support our analysis. This involved converting historical transactions to use the most recent version of the Harmonised System for product classification (HS 2022), establishing firm ownership structures to reconcile subsidiaries at the country level, and addressing the set of firms that do not complete a transformation on the goods they import, or their imports are unrelated to their exports.
HS 2022
As some transactions in our data use the previous 2017 Harmonised System of product classifications we convert these to the 2022 Harmonised System so that we have a consistent set. This was necessary as part of the period under study was in 2021 where the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations. We followed guidance produced by the Statistical Division of the UN for the conversion of products between the different harmonised systems which applies a matching process based on correlation tables.
Firm ownership
Within the data, transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational).
The reconciliation of subsidiaries within country is important as the underlying data has instances whereby a subsidiary of a multi-national imports a product which is then used by a separate subsidiary of the same multi-national. In this case the imported goods are transferred between the subsidiaries domestically (under the same owner umbrella), but we do not capture these transfers as our data contains only cross-border trade.
Using GSCIP ownership data, we complete a process to agglomerate all firms to an ’owner-country’ level.
To rationalise the inconsistent ownership hierarchy, we use GSCIP’s ownership hierarchy dataset which sets out (where available) the global ultimate owner of subsidiaries. For each firm-entity we climb the ownership hierarchical structure until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e. Heinz and Duracell into a larger non-distinct owner such as Berkshire and Hathaway.) To do this we stop climbing the ownership hierarchy to a higher-level firm when that firm can be classified as a financial corporation.
We also faced the issue of circular ownership - whereby a firm is owned by another one or more firms that are (at some tier) owned by the original firm - creating an infinite loop of ownership. To address this problem we break the circular ownership choosing one of the firms within the loop.
The ownership information we rely on is non-uniform and relationships can be opaque. This process improved the information we have on the firm set - but is however not a perfect process.
Once we have assigned subsidiaries to larger conglomerates we create a country-based version of each firm. So that, for example, a multinational firm that transacts to and/or from the UK would be treated as distinct to the same multinational that transacts to and/or from the US.
Additional analysis
Figure S1: [A-B] In and Out Degree. [C] Capital, intermediate and consumption goods (UN definition).
Community detection. We present the number of communities, the Stability coefficient and the Variation of Information for the SC and AI networks for each partition in Figure S2. As the partition number increases, the algorithm runs for a longer ’time’, and larger (but typically fewer) communities are detected.
At each scale (partition number), the algorithm seeks to minimise the Stability coefficient and hence the communities are better defined for lower values. In addition, a low value of the Variation of Information tells us we find similar partitions for different random seeds. We observe that both the SC and AI network have very well defined communities at large partition numbers (low Stability and low Variation of Information), and hence we focus on communities in this region. In Figure S2 C-E, we visualise partitions 20, 25 and 30.
In Figure S3 we show two partitions of the NAFTA IO network, which exhibit similarities to the SC and AI networks.
Figure S2: Community detection statistics: Number of communities, stability and variation of information for SC and AI networks.
Figure S3: IO communities: Two partitions of the Nafta IO network, which exhibit similarities to the SC and AI networks.
10 year growth regressions. In Figure S4 we replicate Figure 7 C and F for 10 year growth regressions (2011-2021). We find similar patterns with a slightly lower AUC in both cases.
Figure S4: 10 year growth regressions:

### v3

@@
 Critical for policy-making and business operations, the study of global supply chains has been severely hampered by a lack of detailed data.
-Here we harness global firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products.
+Here we harness international firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products.
 Transforming this data to a directed network, we find that products are clustered into three large groups including textiles, chemicals and food, and machinery and metals.
-European industrial nations and China dominate critical intermediate products in the network such as metals, common components and tools, while industrial complexity is correlated with embeddedness in densely connected supply chains.
-To validate the network, we find structural similarities with two alternative product networks, one generated via LLM queries and the other derived by NAFTA to track product origins.
-We further detect linkages between products identified in manually mapped single sector supply chains, including electric vehicle batteries and semi-conductors.
-Finally, metrics derived from network structure capturing both forward and backward linkages are able to predict country-product diversification patterns with high accuracy.
-Keywords: Supply chains, resilience, policy, network science
-Disclaimer: This paper contains the views of the authors and not necessarily those of the UK Government or its Ministers.
-Acknowledgements: We would like to thank Christian Diem and Tobias Reisch for their useful input and comments, and Shaun Hoang for his research assistance.
-The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP).
-NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme.
+European industrial nations and China dominate critical intermediate products such as metals, common components and tools, while industrial complexity is highly correlated with embeddedness in densely connected supply chains.
+We find structural similarities with AIPNET, a product network generated via LLM queries, and strong linkages between products identified in manually-mapped electric vehicle battery and semiconductor supply chains.
+Finally, both forward and backward linkages are predictive of country-product diversification patterns, with stronger overall evidence for backward (upstream) linkages.
+Teaser: Harnessing 1 billion cross-border inter-firm transactions to reconstruct the global supply chain production network.
+Disclaimer: Any views expressed are solely those of the author(s) and so cannot be taken to represent those of the Foreign Commonwealth and Development Office (FCDO), or to state FCDO Policy.
+This paper should therefore not be reported as representing the views of the FCDO, its Ministers or His Majesty’s Government.
 Introduction
-As acutely highlighted by the global pandemic, limited insight into global supply chains can have serious and immediate consequences.
+As acutely highlighted by the global pandemic, limited insight into global supply chain dynamics can have serious and immediate consequences.
 However, until quite recently, supply chain analysis was limited to specific sectors or highly aggregate economy-wide transaction data which lacks detail and hence is not very useful or sufficiently accurate for many policy and research questions.
-Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product level supply chain network.
-Traditionally, flows and transactions between sectors in the economy have been captured via input-output tables (Leontief, 1986).
+Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product-level global supply chain network.
+Traditionally, flows and transactions between sectors in the economy have been captured via input-output (IO) tables [1].
 These capture monetary flows between firms within defined industries, and are typically produced by national statistical agencies.
 Some countries, such as the US and Mexico, produce these at quite a detailed level (4 digit NAICS - or approx 400 industries).
-However, in most countries these are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors).
+However, in most countries they are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors).
 And, importantly, they don’t capture cross border flows limiting their potential for global supply chain analysis.
 There have been a handful of efforts to build integrated input-output networks that include inter-country flows, but again these are extremely aggregate in nature.
-For example, the World Input-Output Dataset (WIOD) (Timmer et al., 2012) has just 56 sectors and 40 countries (2010-14) or the OECD World Input-Output Table (WIOT) with 36 sectors and 69 countries (1995-2020).
-A variety of studies have investigated the structure of the World Input-Output Network (WION), finding strong geographical clustering of sectors within neighbouring countries, but also some instances of strong inter-country clustering in sectors such as textiles and transport equipment Piccardi et al.
-(2018).
-In recent years there has been a huge push to better collect and make available for research firm level transaction data (Pichler et al., 2023).
-Research using this data has indicated that, for example, the propagation of shocks at the firm level has distinct characteristics and magnitudes which cannot be approximated by aggregate data (Diem et al., 2022, 2024; Inoue and Todo, 2024).
-Up until very recently, however, most firm level data was collected by single countries (e.g.
-VAT data) or for a specific purpose (i.e., shipping data), and hence had very limited coverage.
+For example, the World Input-Output Dataset (WIOD) [2] has just 56 sectors and 40 countries (2010-14) and the OECD World Input-Output Table (WIOT) has just 36 sectors and 69 countries (1995-2020).
+In recent years there has been a huge push to better collect (and make available for research) firm-level transaction data which has until now been mainly available for individual countries [3].
+Research has found common features across these national datasets [4], and that systemic risk in production networks is concentrated in firms that occupy critical positions in product-level supply chains and provide essential intermediate goods [5, 6, 7].
+This data is typically, however, collected by single countries (e.g.
+VAT data) or for a specific purpose (i.e., shipping data), and hence has very limited coverage when it comes to studying global supply chains.
 At the other end of the spectrum are efforts to manually reconstruct detailed supply chains for specific sectors.
 These studies collect data from a variety of sources including internal systems like procurement, inventory management, and logistics, as well as external sources such as supplier data and transportation information, to build a detailed picture of the key inputs needed to produce often strategically important goods.
-For example, there have been careful efforts to map out antibiotic (Klimek et al., 2023), semi-conductor (Haramboure et al., 2023) and battery supply chains (Caravella et al., 2024).
-However, these supply chain mapping exercises tend to focus on material inputs - critical minerals and chemicals - rather than machinery, tools and other necessary inputs such as cleaning equipment.
+For example, there have been careful efforts to map out antibiotic [8], semi-conductor [9] and critical minerals in electric vehicle battery [10] and defence [11] supply chains.
+However, these mapping efforts tend to focus on material inputs such as critical minerals and chemicals rather than machinery, tools and other necessary inputs such as cleaning equipment.
 And, critically, their fragmented nature makes them not suited to many questions which require a wider structural lens on product inter-connectivity.
-A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks.
-Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters, sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad based shocks such as the pandemic or financial crisis.
+A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks [12].
+Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters [13, 14], sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad-based shocks such as the pandemic [15] or financial crisis.
 Much of this literature aims to quantify node importance - whether it be nodes that influence the degree of spreading processes, or those that are most susceptible to impacts.
-A number of papers suggest that the US IO network follows a power law structure (Acemoglu et al., 2012; Han and Goetz, 2019; Chopra and Khanna, 2015), and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades.
-Blöchl et al.
-(2011) proposed two centrality measures to capture the impact of shocks: random walk centrality to capture most immediate impacts, and betweenness to capture where a shock lingers longest.
-The authors apply these metrics to the OECD WIOT, finding ’wholesale and retail trade’ is most often the node with highest random walk centrality while ’motor vehicles’ have the highest betweenness.
-Supply chain linkages are also key to longer term impacts such as diversification of countries and regions into new economic activities, otherwise referred to as ’upgrading’ (Kaplinsky, 2000).
-Originally proposed as a mechanism to aid countries with natural resources diversify into productive activities (Hirschman, 1958; Gereffi et al., 1999), it is widely understood today that embeddedness in global supply chains aids both domestic productivity and export diversification (Pahl and Timmer, 2020).
-While a wide array of mechanisms have been proposed, many focus on accumulation of technological capabilities and the diffusion of knowhow, as espoused by the Evolutionary Economic Geography (Nelson and Winter, 2002; Boschma and Frenken, 2006) and Innovation Systems (Edquist, 2013) literatures.
-From this perspective, supply chain linkages are thought to drive knowledge diffusion via, for example, demands for quality upgrading, worker flows and increased competition (Blomström and Kokko, 1998; Criscuolo et al., 2017).
-These effects are thought to be particularly strong in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals (Javorcik, 2004), but there are greater uncertainties about the potential for such linkages to induce upgrading when it comes to linkages spanning large geographical and technological distances (Pietrobelli and Rabellotti, 2011).
-The empirical evidence on the benefits of participation in global value chains is mixed (as reviewed by Kowalski et al.
-(2015)), finding a strong influence on manufacturing productivity and export sophistication, but mixed impact, for example, on domestic employment (Pahl and Timmer, 2020; Gonzalez, 2016).
-In terms of direction, both backward linkages (i.e., diversification into more complex downstream goods) and forward linkages (i.e., into input or supplier goods) have been proposed as channels for diversification.
-Methodologically closest to our work, Bahar et al.
-(2019) find evidence for the role of forward linkages in international export product diversification patterns over a 10 year period.
-Linkages in this case, however, were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
+For example, a number of papers suggest that the US sectoral IO network follows a power law structure [16, 17, 18], and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades.
+To improve resilience, some recent work focuses on mitigation strategies such as re-routing [19].
 Beyond the short-medium term impacts of shocks, a recent literature studies the influence of production network structure on local economic growth and resilience.
-Taking an evolutionary stance, in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms (Boschma, 2015).
-Empirical evidence at a subnational level supports the idea that regions’ diversification patterns are conditioned on their ’role’ in national production networks.
-For example, US cities tend to enter new industries that are ’related’ to those that already exist (Essletzbichler, 2017), where relatedness is measured in terms of supply chain proximity (Fan and Lang, 2000).
-When it comes to resilience, Han and Goetz (2019) show that weighted centrality metrics derived from the US IO network are associated with employment resilience for US counties during the 2007-2009 recession, while Cainelli et al.
-(2019) find that EU regions with highly integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
+Taking an evolutionary stance [20], in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms [21].
+For example, [17] show that US regions with more central industries in the US IO network experience better employment resilience during the 2007-2009 recession, while [22] find that EU regions with highly localised and integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
+One of the key mechanisms by which supply chains support local economic growth is through sectoral diversification.
+A key pillar of industrial policy, particularly in resource-rich countries such as African nations with critical minerals, industrial ’upgrading’ is seen as a way for countries to diversify into higher value-added activities [23, 24, 25].
+While a wide array of mechanisms have been proposed, many focus on diffusion of knowhow via, for example, demands for quality upgrading, worker flows and increased competition [26, 27].
+While these effects have been long documented in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals [28], there are greater uncertainties about the potential for such linkages to induce upgrading when spanning large geographical and technological distances [29].
+Known as backward diversification, countries also move into ’upstream’ goods such as machinery or components.
+Hirschman (1958) [23] suggested that the combination of a local market for the good and know-how on how it works would drive this type of diversification, particularly in developing countries.
+Upstream goods include both basic inputs and highly complex and high-value goods such as chemicals, electronics and equipment produced by advanced manufacturing economies [30, 31].
+Various empirical studies find evidence of upstream diversification [32, 33].
+Methodologically closest to our work, Bahar et al (2019) [32] focuses on international export product diversification patterns over a 10 year period, but linkages in this case were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
 Here we aim to infer detailed ’product recipes’, or the set of unique inputs needed in the production of a specific product.
-In other words, we aim to capture the input materials and tools used in product that are specific to a particular product, and not common across many products (e.g., packaging).
-Distinct from two recent efforts build a similar product network using firm-level transaction data from Italy (Fessina et al., 2025) and Japan (Inoue and Todo, 2024), we aim to capture global supply chains (as opposed to local supply chains) via global firm purchasing patterns.
-To do this, we harness ’next generation’ global firm-level transaction data via a UK government programme, GSCIP.
-We use an aggregated subset of data from GSCIP for the years 2021-2023, prior to additional augmentation of the transactional data completed in 2025.
-The subset of data that we use is orders of magnitude larger than single country datasets typically in use today, and covers 20m global firms, and 1 billion cross-border transactions.
-The data was assembled from a variety of sources, from customs to firm-owned data, and matched at a firm level.
-To detect the set of inputs required to make a specific product, we compute the excess purchase of inputs (products) for the set of firms that make a specific product relative to how often an average firm purchases the input.
+In other words, we aim to capture the input materials and tools that are specific to a particular product (and hence not common across many products, e.g., packaging).
+Distinct from two recent efforts build a similar product network using firm-level transaction data from specific countries (Italy [34] and Japan [7]), we aim to capture global supply chains.
+To do this, we harness ’next generation’ global inter-firm transaction data via a UK government programme, GSCIP.
+The data, procured under license by GSCIP, was compiled by a commercial entity from a variety of sources, from customs to firm-owned data, and matched at a firm level.
+To detect the set of inputs required to make a specific product, we compute the excess purchase of each input (product) by the set of firms that make that product relative to how often an average firm purchases the input.
 This latter step effectively down-weights common or ubiquitous inputs.
-This approach echoes that of O’Clery et al.
-(2021), in which the authors inferred country-product diversification paths by extracting the set of common products produced in the set of countries in which a specific new product appeared later.
-In this case, one can think of this approach as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate each set of input linkages for each product.
+This approach echoes that of [35] in which the authors inferred typical product diversification paths by extracting the set of common products produced previously by countries which went on later to make a specific product.
+One can think of our approach here as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate the set of input linkages for each product.
 In this way, we capture both inputs or ingredients, and complex tools and packaging.
@@
 Analysis of the meso-structure of this network suggests that textiles’ supply chains are largely isolated from wider manufacturing, while chemicals are typically more inter-connected with food supply chains than machinery and metals.
-European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of Hidalgo and Hausmann (2009)) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
-While there is no ‘ground truth’ data available, we have undertaken significant efforts to validate the network structure.
-We found structural similarities with an alternative AI-generated product network (Fetzer et al., 2024), and a product supply chain dataset created by NAFTA to track product origins (Conconi et al., 2018).
-In addition, we found dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries and semi-conductors.
-Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy, finding evidence for both forward and backward linkages.
+European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of [36]) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
+While there is no ‘ground truth’ data available, we have undertaken significant efforts to compare the network structure to other related efforts.
+We found structural similarities with an alternative AI-generated product network [37], and a product supply chain dataset created by NAFTA to track product origins [38].
+In addition, we find dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries [10] and semiconductors [9].
+Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy.
+We find evidence for both downstream and upstream linkages, but with overall stronger results for diversification into upstream goods.
 Results
-We use global inter-firm transaction data covering over 20m firms and 1 billion transactions.
-Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as shipping data.
-These sources have been integrated using AI-based tools to, for example, disambiguate firm names and match them.
-The data we used represents a static cut which was made available to the research team via a UK government programme (GSCIP), includes firm level attributes for over 20m firms and 1 billion inter-firm transactions between 2021-23.
-All transactions are between firms located in different countries.
-The GSCIP data contains additional firm-level information from other sources, such as firm ownership structures.
-We have a number of firm and transaction attributes, a selection of which is shown in Figure 1 B, including dollar transactions for specific products between firm pairs.
-As shown in Figure 1 C-D, the data has high coverage across developed economies, and those with high trade (exports plus imports) volumes such as China and Vietnam.
-It also has good coverage in some middle income countries such as India and Peru, but with lower coverage overall in less economically active regions such as Africa.
+We use global inter-firm transaction data covering over 20m firms and 1 billion transactions to construct the product network.
+Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as bill of lading (shipping) data.
+These sources have been integrated by a commercial provider using AI-based tools to, for example, disambiguate firm names and match them.
+The data covers transactions between firms located in different countries during the period 2021-23 and was made available to the research team via a UK government programme (GSCIP).
+GSCIP procured the firm transaction data from a commercial entity, and linked it to additional firm-level information from other sources, such as firm ownership structures (see SI for more detail).
+We have a number of firm and transaction attributes, a selection of which are shown in Figure 1 B, including dollar transactions for specific products between firm pairs.
+While our dataset does not include primary-source data for every country, as shown in Figure 1 C-D, the data has high coverage across developed economies including G7 countries, and those with high trade (exports plus imports) volumes such as China and Vietnam.
+It also has good coverage in some middle income countries such as India and Peru, but with lower numbers of firms overall in less economically active regions such as Africa.
+We provide additional information on coverage in the SI.
 Figure 1: We use global inter-firm transaction data covering over 20m firms and 1 billion transactions.
 [A] Firm level transaction data has been compiled from a variety of sources.
-[B] This static cut of GSCIP data includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions.
+[B] This includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions.
 [C-D] The data has very good coverage of firms in developed economies, with lower coverage in less economically active regions such as Africa.
@@
 In our toy example in Figure 2 B, we show that while aluminium panels are purchased very frequently by washing machine firms (relative to all firms), cardboard boxes are often purchased by all types of firms including washing machine firms and hence they are a very common input (and are thus down-weighted in our scheme).
-More precisely, we set stringent definitions to identify firms that both produce and purchase specific products:
+More precisely, we set stringent conditions to identify firms that both produce and purchase specific products:
   • A firm $k$ produces product $j$ if they sell a larger monetary value of $j$ than the average firm that sells product $j$.
@@
 We can also set a minimum for the size of the set $|S^{i}_{j}|$ (the number of firms that both produce $j$ and buy $i$), which we refer to as parameter ’firmcount’.
-In essence, for low values of firmcount, we have few ’pieces of evidence’ for the edge weight (and so for some applications we might want to set a higher firmcount threshold, which would make our network sparser).
-In the analysis below, we often show results for a range of values of both the edge weight threshold and firmcount.
+For low values of firmcount, we have few ’pieces of evidence’ for the edge weight - and so for some applications we might want to set a higher firmcount threshold.
+This would make our network sparser, but we would have higher confidence in the edge weights as they are based on a higher number of signals from the data.
+In the analysis below, we typically show results for a range of thresholds on both the edge weight ($A_{i,j}$) and firmcount ($|S^{i}_{j}|$).
 As shown in Figure 2 B, we pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
@@
 [B] We pick up both inputs or ingredients and complex tools.
-For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
+For example, for beer, we capture everything from hops and malt to containers and bottle stoppers.
 We transform the product input relations into a network, and find three main clusters representing machinery and metals, chemicals and food, and textiles.
-We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$.
+We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$ products.
 In Figure 3 A, we observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal.
-As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold has no major effect.
-In practice, the majority of the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$.
-In many cases, and also in the SI, we also show results when varying these parameters.
+As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold is less sensitive.
+Unless otherwise stated, the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$.
 This network can be visualised using a ’spring’ based visualisation algorithm (Force Atlas in Gephi was used here), as shown in Figure 3 B.
-We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors.
-However, rather than rely on an external classification which bears no relation to the inter-connectivity we observe in our data, we use a multi-scale community detection algorithm to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’.
-These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network.
-These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation (Fortunato, 2010).
-While there are many distinct approaches to community detection suited to different applications (Fortunato, 2010), in our case, we employ a multi-scale algorithm based on random walkers on a network (Delvenne et al., 2010).
-Essentially, as the walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, we detect larger and larger clusters with strong internal connectivity.
-Hence, as ’time’ goes on, the algorithm detects fewer but larger communities.
-We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters, the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity.
-In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
-Here, shown in Figure 3 C-E, we show three network partitions into communities at three distinct ’times’ (see SI for more details and further analysis).
-The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown.
-We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves).
-We observe that over time, food merges with plastics and then chemicals, and both machinery clusters merge.
-Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged signifying their isolation.
-Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
+We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors, but with strong connections between sectors.
+We use a multi-scale community detection algorithm [39] to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’.
+These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network - and can be seen as an alternative classification of products based on supply chain connectivity.
+These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation [40].
 Figure 3: We transform the product input relations into a network.
@@
 [B] The network can also be visualised using a visualisation algorithm (Force Atlas was used here).
-[C-D] We use a multi-scale community detection algorithm to uncover the ’meso’ structure of the network.
+[C-D] We use a multi-scale community detection algorithm [39] to uncover the ’meso’ structure of the network.
 As the algorithm effectively ’forces’ communities to merge into larger clusters, we find that textiles is highly isolated in the sense that it remains unmerged over many scales.
 Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
+While there are many distinct approaches to community detection suited to different applications [40], in our case, we employ a multi-scale algorithm based on random walkers on a network [39].
+As walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, the algorithm detects larger and larger clusters.
+Hence, as ’time’ increases, the algorithm typically detects fewer but larger communities.
+We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters (although this process is not strictly hierarchical), the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity.
+In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
+Here, in Figure 3 C-E, we show three node partitions corresponding to three distinct ’times’ or scales (see SI Figure S3 for more details and further analysis).
+The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown.
+We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves).
+As the algorithm detects larger clusters over longer times, food merges with plastics and then chemicals, and both machinery clusters merge.
+Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged and are thus highly isolated with weak connectivity to the wider network.
+Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
+Figure 4: European industrial nations and China dominate critical intermediate products, while country-level industrial complexity is correlated with embeddedness in densely connected supply chains.
+[A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ are ’chokepoints’ with high betweenness centrality (BC).
+[B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
+[C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures dense ’neighbour of neighbour’ linkages.
+[D] The Product Complexity Index (rank) correlates 0.46 with hub score.
+[E] Economic Complexity Index (ECI) has a high rank correlation of 0.85 with mean hub score.
+Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
 European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains.
-In the SI we show that machinery products including those related to cars and motors tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to be inputs to many products.
-Further, consistent with what we might expect, deploying a UN classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients.
-Betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross.
-Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) paths transits that node.
-In Figure 4 A, we highlight nodes with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders.
+Moving to node-level metrics, shown in SI Figure S2, we find that machinery products, including those related to cars and motors, tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to have high out-degree (i.e., they are inputs to many products).
+Further, consistent with what we might expect, deploying the UN BEC classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients (SI Figure S2).
+From a network structure perspective, betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross.
+Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) transits that node.
+In Figure 4 A, we highlight products with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders.
 Nations which produce these key inputs have an important role in global supply chains, controlling so-called ’choke points’.
-To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) of 113 countries from global trade data (Comtrade 2024).
+To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) for each country-product pair from global trade data (Comtrade 2024).
 In Figure 4 B, we show a global map with countries shaded by their share of top-50 BC products exported with RCA.
 We find European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
-Highly industrialised nations engage in complex production processes via embeddedness global supply chains.
-To capture production complexity, we adopt an approach proposed by Hidalgo and Hausmann (2009) to infer product complexity (PCI) from patterns of geographical production.
-To measure the complexity of the supply chain in the vicinity of a product in our network, we compute the hub score of each node.
+Next we investigate the extent to which highly industrialised nations engage export goods which are highly embedded in densely inter-connected global supply chains.
+To capture production complexity, we adopt an approach proposed by [36] to infer product complexity (PCI) from patterns of geographical production captured in trade data.
+To measure the localised connectivity and embeddedness of product in global supply chains, we compute the hub score of each node [41].
 This is similar to eigencentrality in the sense that it captures ’neighbour of neighbour’ linkages, but is suitable for directed networks such as ours.
 We find that the PCI (obtained for 2023 from the Harvard Atlas of Economic Complexity) has a rank correlation of 0.46 with node hub score.
-In essence, complex products are embedded in complex supply chains.
-At the country level, the country-level Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score.
-Hence, countries with high industrial complexity tend to export products which are highly embedded in densely connected global supply chains.
-Figure 4: European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains.
-[A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ have high betweenness centrality.
-[B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
-[C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures ’neighbour of neighbour’ linkages.
-[D] The Product Complexity Index (rank) correlates 0.46 with hub score.
-[E] Economic Complexity Index (ECI) has a high rank correlation of 0.72 with mean hub score.
-Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
-The supply chain network correlates with other efforts to reconstruct inter-product linkages.
-First, we consider a recent effect to construct a similar type of product network using LLM queries (Fetzer et al., 2024).
-In essence, the authors asked an LLM what are the top input products for each product, and constructed a network based on its answers in a similar fashion to above.
-Figure 5 A shows that the (binary unweighted) AI network is much sparser than our network (see adjacency matrix in inset).
+In essence, complex products are highly embedded in global supply chains.
+At the country level, the Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score.
+Hence, countries with high industrial complexity tend to export products which are highly embedded in densely inter-connected global supply chains.
+Figure 5: The product network correlates with other efforts to construct inter-product links or chains.
+[A] The ’AI network’ of [37] is constructed via LLM queries.
+Nodes with zero in-degree are shown in white.
+[B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food.
+[C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in our network (SC network).
+[D-E] The correlation of the full edge set of AI network and SC network is approx 0.35, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
+The supply chain network correlates with other efforts to reconstruct inter-product supply chain linkages.
+While there are no ’ground truth’ datasets for us to validate against, and concordances between detailed products and aggregate industries are highly noisy limiting utility of comparing to industry-based Input-Output-type datasets, here we compare our network to a recent attempt to construct a similar type of product network using LLM queries [37].
+Specifically, the authors asked an LLM what are the top input products for each product, and constructed a (binary unweighted) network based on its answers.
+Figure 5 A and B show that the AI network is in general much sparser than our network (we keep the same node layout as above).
 Its clear from the distribution of white nodes (those with zero in-degree) that the AI network tends to have much fewer edges in sectors such as chemicals, plastics and food.
-Overall, our network (the SC network) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community.
+Overall, our network (the ’SC’ network, with parameters set as above) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food community, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community.
 We compare communities for the AI and SC network in Figure 5 C.
-We find that the AI and SC product networks have similar meso-scale structure, with the exception of much larger and stronger chemicals and food clusters in the SC network.
+We find that the AI and SC product networks have a similar meso-scale structure, with the exception of much larger and more strongly connected chemicals and food clusters in the SC network.
 Next, we compute a simple correlation of the full edge set of AI network and SC network, finding they are correlated up to approx 0.35 for low values of firmcount (i.e., including uncertain edges).
 The in- and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree.
-In general, these correlations peak for low firmcount in the range 2-4 - i.e., we keep weak edges for a dense network.
+In general, these correlations peak for low firmcount values in the range 2-4 - i.e., we keep weak edges for a dense network.
 Hence, while the networks look visually very different, they display non-trivial similarity in both meso structure and in- and out-degree pattern.
-We also find similarity to a NAFTA derived product-level IO network.
-We compare our SC network to a NAFTA derived product-level ’IO network’ from Conconi et al.
-(2018) which was constructed based on NAFTA ’Rules of Origin’ which aims to identify the origin country of goods in order to determine if tariffs apply.
+We also compare our SC network to a product network constructed based on NAFTA ’Rules of Origin’ [38] which aims to identify the origin country of goods in order to determine if tariffs apply.
 In this case we find a much lower edge correlation of 0.15, a binary in-degree correlation of 0.28 and out-degree correlation of 0.32.
-Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see Figure S3).
-Figure 5: The product network correlates with other efforts to construct inter-product links or chains.
-[A] The ’AI network’ of Fetzer et al.
-(2024) is constructed via LLM queries.
-Nodes with zero in-degree are shown in white.
-[B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food.
-[C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in the SC network.
-[D-E] The correlation of the full edge set of AI network and SC network is approx 0.34, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
-The product network picks up links identified in manually constructed single-sector supply chains.
-Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector specific supply chains.
-Here we take a selection of these, and investigate to what extent we pick up similar linkages using our method.
-We focus on two key sectors, electric vehicle batteries and semiconductors.
-For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 1 A/C.
-In other words, we show the edge pattern that we find using our method for the identified product sets.
-We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains.
-While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery.
-We wish to show that statistically its very unlikely we could have found these links at random.
-To do this, we simulate 100k random networks with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random networks.
-The modularity score essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard null model which randomly allocates edges while keeping the node degree fixed).
-We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k subgraphs, finding an extremely low probability that our subgraph would be simulated at random (pvalue shown).
-Figure 6: The product network picks up links identified in manually constructed single-sector supply chains.
+Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see SI Figure S4).
+Figure 6: The product network has dense connections between products identified in manually constructed single-sector supply chains.
 [A/D] We capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains.
 [B/E] While products in electric vehicle battery are highly clustered in the centre, semi-conductors exhibit two distinct clusters composed of chemicals and machinery.
-[C/F] To show we pick up linkages between these products that are far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks.
+[C/F] To show that we pick up a pattern of dense linkages between these products which is far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks.
 We observe that our subgraphs are in the very extremes of the distributions suggesting they would be extremely unlikely to occur at random.
-Both downstream and upstream links predict country-product export diversification patterns.
+The product network has dense connections between products identified in manually constructed single-sector supply chains.
+Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector-specific supply chains.
+Here we take a selection of these, and investigate to what extent we pick up similar patterns using our method.
+We focus on two key sectors, electric vehicle batteries [10] and semiconductors [9].
+For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 6 A/D.
+In other words, we show the edge pattern that we find in our network for the manually identified product sets.
+We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains.
+While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery products.
+Could these links be random?
+In other words, if we selected a random set of products, would they also have a similar link density?
+To test this, we simulate 100k random subgraphs with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random subgraphs.
+The (directed) modularity score [42] essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard configuration null model which randomly allocates edges while keeping the node degree fixed).
+We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k random subgraphs, finding an extremely low probability that our subgraph would be simulated at random (p-value shown).
+Figure 7: Upstream links predict country-product export diversification patterns for countries with higher industrial complexity.
+[A/E] Both downstream- and upstream-density captures a country’s presence in products that are downstream/upstream of the target product.
+[B-D] The downstream model (no FE) performs well, in particular for low firmcount (many edges) and top 100 ECI countries.
+[F-H] The upstream model (no FE) performs consistently very well over a wide parameter space, with highest values for low firmcount (many edges), low increase in Rpop, and top 100 ECI countries.
+[I/J] We show the full regression table (include FE) for both downstream and upstream models for high ECI countries and parameters firmcount equals 1 and entry Rpop threshold 0.1.
+Upstream links strongly predict country-product export diversification patterns for countries with higher industrial complexity.
 Finally, we probe whether metrics derived from our network are predictive of country-product diversification patterns.
-To do this, we use a similar approach to a large literature on density metrics in EEG (Hidalgo et al., 2018) to construct a ’downstream’ (and analogous ’upstream’) density metric.
-As illustrated in Figure 7 A/D, essentially, for a product $p$, this captures the share of top in-degree neighbouring products present in country $c$:
+To do this, we use a similar approach to a large literature on density and ’relatedness’ metrics in the Economic Complexity and Evolutionary Economic Geography fields [43].
+Specifically, we construct a ’downstream’ (and analogous ’upstream’) density metric, as illustrated in Figure 7 A/E, for a product $p$ in country $c$ which captures the share of top in-degree neighbouring products of $p$ present in country $c$:
 $d_{p,c}=\frac{\sum_{j\in J_{p}}I(A_{p,j})M_{p,c}}{\sum_{j\in J_{p}}I(A_{p,j})}$
-where $J_{p}$ is the set of top $k=50$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0). We construct an equivalent ’upstream density’ metric.
-Similar to O’Clery et al. (2021), we perform a standard Probit regression for the probability of a product appearance of the form:
+where $J_{p}$ is the set of top $k=40$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0).
+We also construct an equivalent ’upstream density’ metric which considers upstream neighbouring nodes.
+Similar to [35], we perform a standard Probit regression for the probability of a product appearance of the form:
 $R_{p,c}=\Phi(\alpha+\beta_{d}d_{p,c}+\gamma_{p}+\eta_{c})$
 where $R_{p,c}=1$ if product $p$ was present in country $c$ in 2021 but absent in 2016, $\Phi$ is a normal cumulative distribution function, and $\gamma_{p}$ and $\eta_{c}$ are product and country fixed effects respectively.
-The tables in Figure B/E show that both downstream and upstream links predict country-product export diversification patterns with high predictive power (the AUC or ’area under the curve’ measures the model’s ability to distinguish between positive and negative cases) for both models with and without fixed effects.
-The AUC values we find are in line with the literature on predicting product entries in countries, and other similar efforts to predict diversification on a range of scales using density metrics derived from a variety of networks capturing ’capability overlap’ between products (Hidalgo et al., 2007; O’Clery et al., 2021).
-Figure C/F shows, perhaps counter-intuitively but in line with some of the validation steps above, that the predictive power is highest for low firmcount.
+Rpop, which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product, is a threshold used to define product ’presence’ or ’absence’ (similar to Revealed Comparative Advantage [44]).
+In Figure 7 B-D and F-H, we show the predictive power (the AUC or ’area under the curve’) for both models (downsteam and upstream respectively) without fixed effects for three distinct country sets: a full set of 235 countries, the top 100 countries by ECI and the bottom 100 countries by ECI.
+The grids vary parameters firmcount and Rpop.
+Figure 7 J and I show the full econometric model (with various fixed effect models) in table format for the subset of 100 high ECI countries with network parameters firmcount equals 1 and entry Rpop threshold 0.1.
+Overall, across variation in model parameters and country subsets, we find that both downstream and upstream linkages are predictive of diversification patterns - but that upstream linkages are significantly more predictive especially for the higher ECI group of countries.
+This is consistent with an established literature on the role of upstream linkages in diversification processes [23, 32, 33].
+The high ECI set is overall better predicted relative to the full set and low ECI set, which is likely due to the fact that, as we saw above, these countries are more strongly embedded in densely connected global value chains.
+It is also clear that, perhaps counter-intuitively, the predictive power is highest for low firmcount values.
 Hence, including ’weak links’ for which there is relatively lower certainty is informative for the model.
-The predictive power is also higher for lower entry threshold - i.e., we better predict small increases in Rpop (which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product) over a 5 year period compared to larger jumps.
-Figure 7: Both downstream and upstream links predict country-product export diversification patterns.
-[A/D] Downstream/upstream-density captures a country’s presence in products that are downstream/upstream of the target product.
-[B/E] Both downstream and upstream links predict country-product export diversification patterns with fairly high predictive power (AUC) for cases with and without fixed effects.
-[C/F] The predictive power is highest for low firmcount (i.e., including ’weak links’), and for low rpop entry threshold (i.e., we predict small increases in rpop (product concentration relative to GDPPC).
+We are also better in general at predicting small increases in Rpop over a 5 year period compared to larger jumps in Rpop.
+However, many model specifications do achieve high AUC for larger values of Rpop, e.g., for Rpop=0.5, the upstream model (for high ECI countries/low firmcount) has AUC close to 0.8.
+The AUC values we find are in line with (or exceed) the literature on predicting product entries in countries, and other similar efforts to predict diversification using density metrics derived from a variety of networks capturing ’capability overlap’ or linkages between products [45, 35, 46].
+In the SI we show similar results for a 10 year model (2011-21), and variation in density parameter $k$ (this peaks around $k=40$ for most specifications, which is consistent with previous work [35, 46], and used in Figure 7 here).
 Discussion
-In this paper we develop a highly granular network tool to describe complex global inter-related product supply chains.
-We explore its broader structural characteristics, and investigate which countries dominate common input nodes with high betweenness, and which are active in highly embedded ’hub’ nodes.
-Despite an absence of ’ground truth’ data, we validate it in several ways including comparing it to a variety of other efforts to construct product level supply chain networks, and manually constructed data on sector specific supply chains.
+In this paper we develop a highly granular network tool to describe complex global product supply chains.
+We explore its broader structural characteristics, and investigate which countries dominate chokepoint nodes with high betweenness, and which are embedded in highly connected supply chains.
+Despite an absence of ’ground truth’ data, we compare the network to a variety of other efforts to construct product-level supply chain networks, as well as manually constructed sector-specific supply chains.
 There are a wide range of potential research and policy uses for this type of tool.
-From a policy viewpoint, this tool could be used to, for example, quickly map supply chains for new products which have not already been mapped manually, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs.
-More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains (Baldwin, 2006).
+From a policy viewpoint, this tool could be used to, for example, gain an overview of the location of individual products in global supply chains in the absence of manual mapping, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs.
+More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains [47].
+These dynamics have implications for economic inter-dependence between countries, and the increase the potential for economic tools like tariffs and subsidies to exert pressure and influence [48].
 The network structure provides key insights on ’central nodes’ and clusters of highly interconnected products.
@@
 For example, the fact that steel has very high betweenness centrality indicates that it likely plays a role in many paths (chains) and might be seen as a critical good.
-This accords with the fact that the UK government moved aggressively to protect domestic production of domestic virgin steel in early 2025 in the face of impending closure of British Steel.
-More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks and policies such as tariffs in a more detailed, comprehensive manner than previously possible.
-In the case of product clusters, we can see network communities as a sort of alternative product hierarchy relative to official sectoral classifications.
-Highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be collectively targetted by policy interventions rather than individually.
-In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods (Sato, 2014).
-Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions (Hertwich and Wood, 2018).
-This type of detailed network which captures - at least in theory - the full set of product relations could be used to track ’product life-cycle’ emissions as products are manufactured and then used in other manufacturing processes (Wang and Gupta, 2011).
-Here we show that country-product export diversification is associated with both forward and backward linkages, with a slightly stronger signal for backward linkages.
-The network could be used as a policy tool to identify opportunities in the form of upstream and downstream products which could be targets for industrial policy support.
-This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors (Balland et al., 2019).
-Beyond nodes, it also sheds light on potential new edges.
-For example, it could be used to identify linkages which exist globally but not locally.
+This accords with the fact that the US government moved aggressively to protect domestic production of steel throughout 2025 via steep tariffs on imports of both steel and steel ’derivatives’ which includes products with a high steel content such as bolts and bicycles.
+More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks - ranging from financial shocks to conflict to natural disaster - and the impact of policies such as tariffs in a more detailed, comprehensive manner than previously possible.
+For example, highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be jointly targetted by policy interventions rather than individually.
+In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods [49].
+Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions [50].
+This type of detailed network which captures global production paths could be used to track cross-border ’product life-cycle’ emissions as products are manufactured and then exported to be used in other manufacturing processes [51].
+Here we show that country-product export diversification is associated with both forward and backward linkages, with a stronger signal for forward (upstream) linkages, and more complex economies.
+Consistent with the view that upstream products, which are inputs to many other goods, are often highly complex and high value-add, such as semi-conductors or optical microscopes, it seems that recent diversification patterns tend to move in this direction.
+This is consistent with previous findings [32].
+In a similar vein to [45, 22], the network could be used as a policy tool to identify ’feasible’ opportunities in the form of both upstream and downstream products which could be targets for industrial policy support.
+This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors [52].
 There are some limitations to our approach, as well as potential sources of noise.
-In order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years.
-This approach, however, does not currently allow for temporal analysis.
+Firstly, in order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years.
+This approach, however, does not allow for temporal analysis.
 Future work may investigate the extent to which disaggregating into distinct time periods and recomputing the network for each is feasible.
@@
 Third, there is no doubt some noise in our data due to the uneven composition of the firm-level data.
-For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of firms in our dataset, and some countries/sectors will be more represented than others.
+For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of all global firms in our dataset, and some countries/sectors will be more represented than others.
 This issue is less problematic for our tool than say country-specific or firm-level analysis, but nonetheless it will introduce noise which we cannot fully mitigate against.
@@
 For example, we noticed an unusual input to semiconductors - washing machines.
-But it turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips.
+It turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips [53].
 Another source of noise is likely to be multi-product firms.
-While we have taken efforts to reduce this, by omitting firms that sell products in more than five HS Sections, and looking for common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
+While we have taken efforts to reduce this by omitting firms that sell products in more than five HS Sections, and our overall approach based on common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
 Fifth, there may be idiosyncratic differences between supply chain relationships within regions that we aggregate over by computing a network using all firm transactions.
 For example, battery supply chains may look different in the US vs Asia.
-While we have looked at some geographic variations here (see SI), in future work we may reconstruct the network using data for specific regions to further investigate this issue.
+In future work we may reconstruct the network using data for specific regions to further investigate this issue.
 Finally, we present the network at HS 4-digit level.
-It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and a more complicated validation exercise with fewer sources of external data to validate against.
+It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and fewer sources of external data to compare the network structure to.
 Again, we leave this to future work.
-Methods
+Materials and Methods
 Processing of underlying firm data
-The cut of data we used from GSCIP on firm-level transactions covers the time period between Jan 2021 and Dec 2023.
-We take the following processing steps:
+Our GSCIP data includes firm-level transactions during the time period between January 2021 and December 2023. We take the following processing steps:
   • We assign all firms to an ’owner firm’ based on ownership information available to us (see SI).
-  • We aggregate all firms to an ’owner-country’ level. In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
-Construction of supply chain network
+  • We aggregate all firms to an ’owner-country’ level (see SI). In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
+Multi-product firms
 A concern about our approach might be that it could pick up spurious links due to the presence of firms in our dataset which make many products (and hence we cannot infer specific inputs to products from their purchases).
 Our methodology is designed to reduce this issue since we look at the purchases of not just one firm but all firms that make a specific product.
-To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of products such as Proctor and Gamble).
+To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of diverse products such as Proctor and Gamble).
 In order to further reduce noise, in addition to the firmcount parameter as mentioned above, we also drop any links from product $i$ to product $j$ in which the total value of the transactions of the firms in set $S^{i}_{j}$ is less than \$1000.
 Network analysis
-Community detection We use the Stability algorithm of Delvenne et al.
-(2010) to detect communities in the SC and AI networks.
-The algorithm detects partitions on a range of scales corresponding to random walker time on the network.
-The higher the resolution, the longer the time, the larger (and fewer) the communities.
+Community detection We use the Stability algorithm of [39] to detect communities in the SC and AI networks.
+The algorithm detects partitions on a range of ’scales’ corresponding to the ’time’ spent by a random walker exploring the network.
+The lower the time, the smaller (but more numerous) the communities detected.
+The longer the time, the larger (and fewer) the communities.
 We use the ’full’ (non-linearised) specification for a directed Laplacian, with 100 iterations.
@@
 Modularity score We use the Modularity score to compare the connectivity of the subgraph induced by each of our sector-specific supply chains (EV batteries and semi-conductors) to 100k randomly generated subgraphs of the same size.
-Modularity (Newman, 2006) compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
-Effectively, we decompose the directed Modularity score proposed by Leicht and Newman (2008) to evaluate a set of communities for a full network, and use it instead to evaluate the connectivity of a subgraph (i.e., a single community).
+Modularity [54] compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
+Effectively, we decompose the directed Modularity score proposed by [42], which aims to evaluate the connectivity of a set of communities in a network, and use it instead to evaluate the connectivity of a single subgraph (i.e., a single community).
 This is mathematically coherent as Modularity is a sum over the connectivity within communities.
@@
 To capture product entries, we wish to compute a measure of a countries’ concentration in a product.
-The traditional approach is Revealed Comparative Advantage (RCA) proposed by Balassa (1965) captures is the proportion of the country’s exports in a product divided by the global proportion of exports in a product.
-We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates (Hidalgo, 2021).
+The traditional approach is Revealed Comparative Advantage (RCA), proposed by [44], captures the proportion of the country’s exports in a product divided by the global proportion of exports in a product.
+We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates [55].
 Specifically, it compares a country’s per capita production levels in a product to the world’s overall per capita production of the product:
@@
 where $E_{p,c}$ is the total exports of country $c$ in product $p$, $E_{p}$ is global exports of product $p$, $pop_{c}$ is the total population of country $c$ and poppop is the global total population.
-We conduct the econometric analysis for a set of 235 countries.
-We drop products not in the main connected component of the network and those with relatively small global trade (less than 2 billion\$ in 2016) for a total of 683 products.
-Of a possible 235x683 observations, 86% were absent ($Rpop<0.05$) in 2016.
-Of these, 3.5% of products appeared by 2021 (with $Rpop>0.1$).
-While we keep the absence threshold fixed, we show results for a range of presence thresholds in Figure 7 C and F.
-Supplementary Information
-Overview of underlying data
+We conduct the full econometric analysis for a set of 235 countries.
+We drop products not in the main connected component of the network, which is dependent on the values of parameters firmcount and edge threshold, and 28 products with global trade in 2021 less than \$10m.
+For the tables shown in Figure 7 I/J, with both firmcount and edge threshold equal to 1, we have $n=1198$.
+Of a possible 100x1198 observations in Figure 7 I (high ECI country set), 76% were absent ($Rpop<0.05$) in 2016.
+Of these, 3.6% of products appeared by 2021 (with $Rpop>0.1$).
+Acknowledgements
+We would like to thank Christian Diem, Tobias Reisch, Adnan Khan and Fergus Cumming for their useful input and comments, and Shaun Hoang for his research assistance.
+The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP).
+Funding: NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme (ES/Y004817/1_583688).
+Author contributions:
+Conceptualization: NOC, DTH
+Methodology: NOC, BRB, TS, DTH
+Investigation: NOC, BRB, TS
+Visualization: NOC
+Supervision: NOC, BRB, DTH
+Writing-original draft: NOC
+Writing-review & editing: NOC, BRB, TS, DTH
+Competing interests: Any views expressed are solely those of the author(s) and so cannot be taken to represent those of the Foreign Commonwealth and Development Office (FCDO), or to state FCDO Policy.
+This paper should therefore not be reported as representing the views of the FCDO, its Ministers or His Majesty’s Government.
+Data and materials availability: The underlying firm-level data was made available to the research team via the UK Government Global Supply Chain Intelligence Programme (GSCIP).
+This sensitive commercial data is held under license is, unfortunately, not available to external researchers.
+Supplementary Materials for
+Deciphering the global production network from cross-border firm transactions
+Neave O’Clery* et al.
+*Corresponding author. Email: n.oclery@ucl.ac.uk
+This PDF file includes:
+Supplementary Text
+Figs. S1 to S6
+Supplementary Text
+Data
 Data sources and coverage
-This paper relies on data provided by the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP).
-This contains firm-to-firm cross-border trade, derived from customs and shipping data, with details on the transacting firms compiled from tax filings, company websites, and other sources.
-Firm-to-firm transaction data is derived from administrative customs data.
-GSCIP amalgamates first-party customs data across a multitude of countries.
-Our dataset includes transaction information for every country.
-While the dataset does not have first-hand customs data or another primary source of transaction information for every country, the data that we do have includes transactions between firms within the countries included and those out of sample providing partial coverage.
-We are using a static cut of the data, which has further been restricted to the period from January 2021 to December 2023 as this period had the greatest coverage for the subset.
-Data processing of underlying firm data
-Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional processing of the data to support our analysis.
-This involved converting historical transactions to use the most recent version of the Harmonised System for product classification (HS 2022), establishing firm ownership structures to reconcile subsidiaries at the country level, and addressing the set of firms that do not complete a transformation on the goods they import, or their imports are unrelated to their exports.
+This paper relies on data accessed through the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP).
+This contains firm-to-firm cross-border transactions, derived from customs and shipping data, which was compiled by a commercial provider.
+Further information on firms in the transaction dataset has been gathered by GSCIP from business registries, tax filings, company websites, and other sources.
+The dataset includes ’primary’ customs data from a set of available country sources.
+All other countries are represented in the dataset as both customers and sellers of these primary countries.
+The exact set of primary countries included within the GSCIP programme is suppressed for reasons of sensitivity.
+We do, however, provide an overview of country-level coverage relative to total trade volumes in Figure 1.
+While GSCIP is an ongoing programme, we use data from the period from January 2021 to December 2023, accessed in early 2025.
+Processing of underlying firm data
+Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional cleaning and processing of the raw data as detailed below.
 HS 2022
 As some transactions in our data use the previous 2017 Harmonised System of product classifications we convert these to the 2022 Harmonised System so that we have a consistent set.
-This was necessary as part of the period under study was in 2021 where the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations.
-We followed guidance produced by the Statistical Division of the UN for the conversion of products between the different harmonised systems which applies a matching process based on correlation tables.
+This was necessary as our data includes 2021 when the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations.
+We followed concordance tables produced by the Statistical Division of the UN [56] for the conversion of products between the different harmonised systems.
 Firm ownership
-Within the data, transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational).
+Using GSCIP ownership data, we complete a process to agglomerate all firms to an ’owner-country’ level.
 The reconciliation of subsidiaries within country is important as the underlying data has instances whereby a subsidiary of a multi-national imports a product which is then used by a separate subsidiary of the same multi-national.
 In this case the imported goods are transferred between the subsidiaries domestically (under the same owner umbrella), but we do not capture these transfers as our data contains only cross-border trade.
-Using GSCIP ownership data, we complete a process to agglomerate all firms to an ’owner-country’ level.
-To rationalise the inconsistent ownership hierarchy, we use GSCIP’s ownership hierarchy dataset which sets out (where available) the global ultimate owner of subsidiaries.
-For each firm-entity we climb the ownership hierarchical structure until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e.
-Heinz and Duracell into a larger non-distinct owner such as Berkshire and Hathaway.) To do this we stop climbing the ownership hierarchy to a higher-level firm when that firm can be classified as a financial corporation.
-We also faced the issue of circular ownership - whereby a firm is owned by another one or more firms that are (at some tier) owned by the original firm - creating an infinite loop of ownership.
-To address this problem we break the circular ownership choosing one of the firms within the loop.
-The ownership information we rely on is non-uniform and relationships can be opaque. This process improved the information we have on the firm set - but is however not a perfect process.
-Once we have assigned subsidiaries to larger conglomerates we create a country-based version of each firm.
-So that, for example, a multinational firm that transacts to and/or from the UK would be treated as distinct to the same multinational that transacts to and/or from the US.
+For example, Cadberrys may import cocoa beans using one subsidiary and produce a chocolate bar using another subsidiary.
+Transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational).
+For each firm-entity we climb the ownership hierarchical structure, provided by GSCIP, until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e.
+Heinz and Duracell into a larger non-distinct owner such as Berkshire Hathaway).
+To do this we stop climbing the ownership hierarchy when the next firm up is a financial entity or corporation.
+Firm-level transaction network
+Here we present descriptive statistics on the underlying inter-firm transaction data.
+Degree statistics
+Similar to other work to understand the structure of firm-level production networks [4], we analyse the in- and out-degree of our firm population.
+The 1 billion transactions in our data correspond to 47 million firm-to-firm links.
+We consider this data to form a network in which nodes are firms and weighted edges correspond to the sum of the value of transactions between these firms within our time-period.
+The network is highly heterogeneous: the majority of firms (63%) maintain trading relationships with just one partner, while a minority engage with a large number of partners.
+We find that the maximum out-degree (number of export partners for a single firm) reaches 351,000, and the maximum in-degree (number of import partners) is 113,000.
+Notably, 1% of all firms have more than 50 trading partners, while only 0.01% have over 1,000 partners.
+Within our full dataset of 47 million firms, we observe that 43% are exclusively exporters, 48% are exclusively importers, and only 9% both import and export.
+This latter set includes a core set of 1.8 million firms which are use to construct our production network (the remainder - exporter- and importer-only firms - are origins or destinations of transactions from this core set).
+Firm types by geography
+There are 8,895,000 firms (43% of all firms in the dataset) that export only (i.e., without any recorded imports).
+This pattern is especially strong in China (90.1% sender-only), Taiwan (81.1%), Switzerland (80.0%), Hong Kong (75.3%), Japan (72.8%), and South Korea (70.6%).
+There are 9,771,000 nodes that only import (have no recorded exports).
+The representation of import-only nodes is most notable in Ecuador (88.8%), Peru (87%), the Philippines (81.5%), Canada (74.7%), and Ukraine (67.6%).
+There are 1,803,000 firms (9% of all firms) that export and import within our dataset.
+These firms are especially prevalent in Vietnam (21%), Turkey (18%), the United Kingdom (17%), India (16%) and Germany (16%).
 Additional analysis
-Figure S1: [A-B] In and Out Degree. [C] Capital, intermediate and consumption goods (UN definition).
-Community detection.
+Formulas
+For our main analysis, we set stringent conditions to identify firms that both produce and purchase specific products:
+  • A firm $k$ produces product $j$ if they sell a larger monetary value of $j$ than the average firm that sells product $j$.
+  • A firm $k$ buys product $i$ if they purchase a larger monetary value of $i$ than the average firm that buys product $i$.
+We also produced alternative networks with variations to these criteria, including:
+  1. 1.
+    filtering the network to just include firms buying more of product $i$ than the average firm, without any conditions on selling product $j$.
+  2. 2.
+    filtering for just firms selling more of product $i$ than the average firm, without any conditions on buying product $i$.
+  3. 3.
+    Including all firms selling product $j$ and buying product $i$.
+We replicate Figure 5 in the main text which displays edge, in-, and out-degree correlation with the ’AI network’ in Figure S1.
+We find that the conditions we apply in the main analysis result in stronger correlations with the AI network, regardless of thresholds on edge or firm count.
+Community detection
 We present the number of communities, the Stability coefficient and the Variation of Information for the SC and AI networks for each partition in Figure S3.
@@
 At each scale (partition number), the algorithm seeks to minimise the Stability coefficient and hence the communities are better defined for lower values.
-In addition, a low value of the Variation of Information tells us we find similar partitions for different random seeds.
+In addition, a low value of the Variation of Information tells us we find similar partitions across a 1000 random seeds.
 We observe that both the SC and AI network have very well defined communities at large partition numbers (low Stability and low Variation of Information), and hence we focus on communities in this region.
 In Figure S3 C-E, we visualise partitions 20, 25 and 30.
-In Figure S3 we show two partitions of the NAFTA IO network, which exhibit similarities to the SC and AI networks.
+In Figure S4 we show two partitions of the product network constructed based on NAFTA ’Rules of Origin’ [38] which exhibit similarities to the SC and AI networks.
+Additional econometric regressions
+In Figure S5 we replicate Figure 7 B-D and F-H for 10 year growth regressions (2011-2021), finding very similar results.
+In Figure S5, we vary the ’number of neighbours’ parameter in the density metric for high/low ECI countries, for both downstream and upstream models.
+Overall there is a small but perceptible peak in the AUF (no FE) around $k=40$ for most models, and so we use this value in the main paper.
+This is consistent with previous work [35] which found a peak of around $k=50$ in similar contexts.
+Figures
+Figure S1: Variations in the underlying formula that forms the product network: edge, in and out-degree correlations with the AI network.
+Figures show variation for setting Firm count and edge threshold at 1, 2, 3, 4 and 5.
+The correlation of the full edge set of AI network and SC network is highest with thresholds for both export and import.
+Figure S2: [A-B] In and Out Degree. [C] Capital, intermediate and consumption goods (UN definition).
 Figure S3: Community detection statistics: Number of communities, stability and variation of information for SC and AI networks.
 Figure S4: IO communities: Two partitions of the NAFTA IO network, which exhibit similarities to the SC and AI networks.
-10 year growth regressions. In Figure S4 we replicate Figure 7 C and F for 10 year growth regressions (2011-2021). We find similar patterns with a slightly lower AUC in both cases.
-Figure S4: 10 year growth regressions:
+Figure S5: 10 year growth regressions: all countries, high ECI countries, low ECI countries.
+Figure S6: Variation in density parameter k for high and low ECI country sets. We see the AUC peaks around k=40 for all cases (more darker squares).